2. You are an international shrimp trader. A food producer in the Czech Republic offers to pay you 2.5 million Czech koruna today in exchange for a​ year's supply of frozen shrimp. Your Thai supplier will provide you with the same supply for 3.1 million Thai baht today. If the current competitive market exchange rates are 25.68 koruna per dollar and 40.99 baht per​ dollar, what is the value of this exchange to​ you? The cost of the shrimp from the Thai supplier in dollars is ​$_____ ​(Round to the nearest​ cent.)
Answer #1

Amount paid to the Thai supplier in dollars(P) = 3.1million/40.99 = $75628.202

Amount received from Czech Republic(R) = 2.5million/25.68 = $97352.0249

Value of this exchange to us = R - P = 97352.0249 - 75628.202

= $21723.8229 = $21,723.82(approximately)

The cost of the shrimp from the Thai supplier in dollars is ​$75628.20